OpenStudy (anonymous):

Can anyone help me solve this? I will give medal to whoever helps! 3 Sqrt. (2^6)^x

OpenStudy (anonymous):

\[4^{x}\] this is because exponents of exponents are the same when multiplied. i.e. \[2^{6^{x}} = 2^{6x}\] and roots become exponent division so \[\sqrt[3]{2^{6^{x}}} = 2^{2^{x}} = 4^{x}\]

OpenStudy (anonymous):

\[\sqrt[3]{\left( 2^{6} \right)}^{x}\]

OpenStudy (anonymous):

okay! thanks! could you explain that second part in a little more detail? The first part makes sense... but what do you do with the 3?

OpenStudy (anonymous):

so after you change it to \[\sqrt[3]{2^{6x}}\] what do you do?

OpenStudy (anonymous):

sure, happy to help! Every root is actually an exponent of a fractional amount. \[\sqrt[2]{x} = x ^{\frac{ 1 }{ 2 }}\] so \[\sqrt[3]{2^{6^{x}}} = (2^{6\times \frac{ 1 }{ 3 }})^{x} = 2^{2^{x}}\]

OpenStudy (anonymous):

oh, okay!! thank you soooo much!! :))

OpenStudy (anonymous):

Hey, np! Fan me and ask me questions as much as you like :-D